# An embedding-based distance for temporal graphs

### v2

@@ Abstract
Keywords: Network science, Temporal networks, Graph distance, Graph embedding
@@ Introduction
Graphs are ubiquitous mathematical entities formed by a set of nodes and one of edges connecting them [1, 2, 3]. They can model a wide range of interacting systems, such as social [4, 5], technological [6], spatial [7] and biological networks [8], and owe their popularity to their ability to encode the complex relational structure of these systems. Real-world systems, moreover, often have temporal properties that cannot be encoded in static graphs and call for modeling based on time-resolved network representations, known as temporal graphs [9]. Examples of these include transportation [10] and ecological networks [11], human close-range interactions [12, 13, 14, 15, 16], collaboration networks [17, 18, 19], etc.
@@
In the case of temporal graphs, the temporal dimension adds to the complexity of graph comparison. Many of the distance measures mentioned above for static graphs were used to study temporal graphs [35], typically to compare the state of the graph at different points in time and to identify anomalies and change points in the temporal evolution of the graph [36, 37, 38, 39, 40]. Here instead, we study the problem of comparing temporal graphs, each taken as a whole. Unlike the snapshot-based methods mentioned above [36, 37, 38, 39, 40], we rely on a temporal embedding with a dimensionality that is independent of the temporal span of the graphs, so that we can compare temporal graphs with arbitrary, and different, temporal spans. Few notions of distance were proposed to compare temporal graphs with arbitrary time spans. In particular, the method of Ref. [41] is based on a set of features that can be tailored to a specific research question (similarly to [29] for static graphs). Ref. [42] on the other hand addresses the same problem of temporal graph comparison we tackle here, but it assumes the graphs to be matched and it is computationally inefficient (the graph dissimilarity is obtained solving an NP-hard optimization problem and the authors propose an approximation running in $\mathcal{O}(n^{3})$ operations, where $n$ is the number of graph nodes). Finally, Ref. [43] only considers the unmatched case and focuses on a specific property of temporal graphs, namely the comparison of temporal paths statistics, to provide a definition of dissimilarity.
Here, we propose a computationally efficient method to compute an actual distance metric between pairs of temporal graphs, considering both the matched and unmatched cases. We build our comparison on top of graph embeddings, leveraging ideas from the local and global approaches mentioned above and encoding differences at the topological and temporal levels. Commented Python code that implements the approach described here is available at github.com/lorenzodallamico/G-DynaDist.
Temporal graphs definitions
@@
A temporal graph $\mathcal{G}(\mathcal{V},\mathcal{E},\mathcal{W})$ is a tuple $(\mathcal{V},\mathcal{E},\mathcal{W})$, where $\mathcal{V}$ is a set of $n$ nodes, $\mathcal{E}$ a set of temporal edges between nodes, and $\mathcal{W}$ a set of edge weights. For a discrete variable $t$ denoting time, all temporal edges $e\in\mathcal{E}$ have the form $e=(i,j,t)$, representing an interaction between nodes $i$ and $j$ at time $t$ and $w_{e}$ indicates the positive weight of edge $e$. We say an edge $(ij)$ is active at time $t$ if $(i,j,t)\in\mathcal{E}$. A node $i$ is active at time $t$ if it belongs to an active edge at $t$.
@@
Figure 1: Distance computation between two temporal graphs. The inputs to the distance function are two temporal graphs $\mathcal{G}_{1}$ (top, orange nodes) and $\mathcal{G}_{2}$ (bottom, green nodes), generally with a different number of snapshots ($T_{1}=5$, $T_{2}=4$, in the example) and a different number of nodes $n$ (here $n_{1}=6$, $n_{2}=5$). Each graph is represented as a matrix $P\in\mathbb{R}^{n\times n}$, according to (1), with entry $(ij)$ encoding the probability that a random walker goes from $i$ to $j$ following a time-respecting path (the walker’s position at each time point is indicated in blue). The matrices $P_{1},P_{2}$ are then embedded using the EDRep algorithm, mapping them to $X_{1}\in\mathbb{R}^{n_{1}\times d}$ and $X_{2}=\mathbb{R}^{n_{2}\times d}$. Finally, the matched – (3) – and unmatched – (4) – distances are calculated. Notice that a necessary condition to compute $d_{\rm m}$ is that $n_{1}=n_{2}$.
Main result
In the following, we define two types of distance functions: $d_{\rm m},d_{\rm u}$ between matched and unmatched graphs, respectively. The inputs to these functions are two temporal graphs, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, with potentially different numbers of time points (snapshots) $T_{1}$ and $T_{2}$. In both cases, the distances should fulfill the following standard requirements [45]: i) non-negativity; ii) separation axiom; iii) symmetry; iv) triangle inequality. In addition, both distances should be able by design to detect differences induced by the permutation of time indices. The matched distance should also differentiate between temporal networks differing only by a permutation of node indices, while the unmatched one should be invariant under such transformation.
Our proposed distances are computed in two steps: i) we generate a temporal graph embedding, given by a matrix of size $(n\times d)$ (with $d$ independent of $T$), that encodes relevant topological and temporal properties by leveraging time-respecting random walks on the temporal network; ii) we define the distance in the embedded space, treating separately the matched and the unmatched cases. The distance computation pipeline is illustrated in Fig. 1. The steps outlined above are described in detail in the next two Sections titled Temporal graph embedding and Defining an embedding-based distance.
@@
Graphs are rich data representations, making it challenging to define simple mathematical operations to manipulate them. Graph embeddings provide an interesting avenue for doing so, as they yield a representation in vector space that preserves some of the graphs’ relevant properties [46, 47, 48, 49]. In the case of temporal graphs, the additional challenge of dealing with their temporal dimension has been tackled in many different ways: Ref. [50] represents time as an additional dimension of an adjacency tensor; in evolutionary spectral clustering [51, 52, 53, 54, 55], a different embedding is obtained for each snapshot, and smoothness conditions are implemented on their temporal evolution; more recent work [56, 57, 58] introduced approaches based on time-respecting random walks, which encode temporal properties relevant for dynamical processes occurring on temporal graphs; Refs. [39, 40] also build one embedding per snapshot for studying the graph evolution and comparing different time points. The authors argue that a graph comparison metric should depend on the whole graph structure and thus define graph instead of node embeddings.
With a similar approach to [58], we start from time-respecting random walks to generate temporal graph embeddings, which we use to define a distance based on global temporal graph properties. The embeddings are obtained as the solution to an optimization problem. We define a loss function whose argument is a transition matrix $P$ that is the limiting distribution of time-respecting random walks over the temporal graph $\mathcal{G}$ to be embedded. $T$ being the number of snapshots of $\mathcal{G}$, we consider an ensemble of random walks of length $\ell$ drawn uniformly at random between $1$ and $T$, and starting from a randomly chosen node at time $t=T-\ell+1$; at each time step, the walker, situated, e.g., on node $i$, can either move to a random neighbor of $i$ in the snapshot at time $t$ or stay in place on $i$; if the walker is situated on a node that is not active at time $t$, i.e. , which has no neighbors in this snapshot, the walker stays in place with probability one. The graph snapshots on the left of Fig. 1 show examples of such time-respecting random walks on the input graphs.
Specifically, given a temporal graph $\mathcal{G}$, with $W_{t}$ the snapshot adjacency matrices at times $t=1,\cdots,T$, we define for each snapshot $\hat{L}_{t}=(D_{t}+I_{n})^{-1}(W_{t}+I_{n})$, where $D_{t}$ is the degree matrix $D_{t}={\rm diag}(W_{t}\bm{1}_{n})$ of the snapshot, $I_{n}$ is the identity matrix of size $n\times n$, and $\bm{1}_{n}$ is the $n$-sized unity vector. We then define the matrix $P$ as
@@
The matrices $\hat{L}_{t}$ and $P$ are sometimes referred to as temporal transition matrix and global transition matrix, respectively. The properties of time-respecting random walks depend on several properties of the temporal graph, that may include, for instance, the presence of a non-Markovian behavior, and of a broad distribution of the edge activity and inter-event durations [59]. The limiting probability matrix $P$ of time respecting random walks encodes the aforementioned properties and it is sensitive to the permutation of time indices. We remark that the random walks we use can be of arbitrary length. Hence, the matrix $P$ is, in principle, sensitive to all time scales of the temporal graph at hand.
Given the above probability matrix $P$ and a chosen embedding dimensionality $d$, we now generate, for each node $i$ of $\mathcal{G}$, a unitary node embedding vector $\bm{x}_{i}\in\mathbb{R}^{d}$. To this end, we define $X\in\mathbb{R}^{n\times d}$ as a matrix where each row represents an embedding vector. We then express the loss function as:
$\mathcal{L}_{P}(X)$ $=-\sum_{i,j\in\mathcal{V}}\left[P_{ij}~{}{\rm log}~{}Q_{ij}(X)-\frac{\bm{x}_{i}^{T}\bm{x}_{j}}{n}\right],$ (2)
$Q_{ij}(X)$ $=\frac{e^{\bm{x}_{i}^{T}\bm{x}_{j}}}{\sum_{k\in\mathcal{V}}e^{\bm{x}_{i}^{T}\bm{x}_{k}}}:=\frac{e^{\bm{x}_{i}^{T}\bm{x}_{j}}}{Z_{i}}~{}.$
@@
We want to define a distance between temporal graph embeddings satisfying the requirements stated above. Notice that the embedding vectors $\bm{x}$ are defined up to an orthogonal transformation applied to the rows of $X$. In fact, suppose $R\in\mathbb{R}^{d\times d}$ is an orthogonal matrix, and $\tilde{\bm{x}}_{i}=R\bm{x}_{i}$, then $\bm{x}_{i}^{T}\bm{x}_{j}=\bm{\tilde{x}}_{i}^{T}\bm{\tilde{x}}_{j}$ and the loss function of (2) is $\mathcal{L}_{P}(\tilde{X})=\mathcal{L}_{P}(X)$. Hence, our distance must be invariant under orthogonal transformations. Moreover, as discussed above, the distance for unmatched graphs must also be invariant with respect to node permutations.
Given these requirements, we now discuss separately our choices for the the matched and unmatched cases. In the Empirical evaluation section, we show the ability of our distances to discriminate temporal graph classes that differ for topological or temporal properties.
Distance for matched graphs
@@
We can easily verify that $d_{\rm m}$ is a distance, as it satisfies all the requirements mentioned above and is generally not invariant under the permutation of node indices. Notice that, even if $d_{\rm m}$ can be computed for any pair of graphs with the same size, this distance is only meaningful for matched graphs (i.e., with known matching). In (3), the term $X_{1}^{T}X_{2}$ implies that $X_{1}$ and $X_{2}$ have the same size and that their rows are ordered according to the same node index.
Distance for unmatched graphs
@@
In the Methods section we show that the covariance matrices $X_{1}^{T}X_{1}$ and $X_{2}^{T}X_{2}$ (of size $d\times d$) are indeed an appropriate argument of the distance function $d_{\rm u}$, since they are independent of the ordering of the nodes and any other orthogonal transformation applied to the columns of $X_{1}$ and $X_{2}$. However, a direct comparison of the entries of $X_{1}^{T}X_{1}$ and $X_{2}^{T}X_{2}$ would require a one-to-one correspondence between the embedding dimensions, i.e., the columns of $X_{1}$ and $X_{2}$. This mapping is unavailable because the EDRep algorithm, like most vector-space embedding techniques, makes no guarantees on the correspondence between the dimensions of the embedding vectors corresponding to different graphs. To tackle this issue, we thus compare the spectra, which are invariant with respect to orthogonal transformations of the rows of the embedding matrices. This choice takes inspiration from spectral distance definitions for static graphs, such as Ref. [33], in which graphs are compared by computing the distance between the ordered sets of the eigenvalues of their matrix representations (e.g., adjacency, Laplacian, non-backtracking, etc.)
@@
As discussed in the Methods section, the computational complexity of the embedding step is $\mathcal{O}\left(nd^{2}+d\cdot{\rm min}(n^{2},E)\right)$, where $E=\sum_{t}|\mathcal{E}_{t}|$ is the total number of temporal edges. On top of this, the distance $d_{\rm m}$ is then obtained with $\mathcal{O}(nd^{2}+d^{2})$ additional operations, while $d_{\rm u}$ requires $\mathcal{O}(nd^{2}+d^{3})$ additional operations.
Empirical evaluation
@@
Figure 2: Validation of the distances on graphs of varying size. Panel A: Accuracy of the distance-based clustering against the ground truth classes in terms of normalized mutual information (NMI) as a function of the embedding dimensionality $d$ used in the embedding step shown in Figure 1. The clustering task consists in recognizing four synthetic temporal graph classes with an unsupervised algorithm based on the distance $d_{\rm u}$. The synthetic graphs classes are obtained by generating a graph from either of four models (stochastic block model (SBM), configuration model (CM), Erdős-Renyi (ER), and geometric model (GM)) with constant degree equal to $4.8$, and the temporal component is obtained by sampling the edge activity of an empirical graph, as detailed in the main text. Inset of panel A: Scatter plot of UMAP dimensionality reduction in two dimensions of the vectors $\bm{\lambda}$ appearing in the definition of $d_{\rm u}$ given in (4), with here $d=32$. Each point refers to a temporal graph; the color and marker code refer to the generative model of its static component, while the marker size is proportional to the number of nodes. Panel B: Each point is the $2$-dimensional UMAP embedding of the vector $\bm{\lambda}$ of the temporal graph obtained selecting two classes and a day of interaction for all possible $(c_{2},c_{2},day)$ triplets in the SocioPatterns data sets describing proximity temporal graphs in schools. The color and marker code are assigned according to the class labels: the primary school classes form a group on their own and the other three groups (Other-Other, Bio-Other, Bio-Bio) belong to the high school data sets, where Bio are the biology classes, and Other the remaining ones. In all cases, we use a temporal resolution equal to $10$ minutes to create the temporal graphs from the SocioPatterns data. Panel C: normalized matched distance between a temporal graph and itself upon partial node re-labeling, as a function of the fraction $\alpha$ of re-labeled nodes. Each plot considers one of the generative models used in panel A, with the same nomenclature used for the title. The – barely visible – shadow line is the standard deviation of the distance value across $25$ randomizations of the partial re-labeling.
@@
Figure 2A shows the normalized mutual information (NMI) between the known temporal graph classes (associated with the four generative models) and the discovered cluster labels, as a function of the embedding dimension $d$. The discrimination performance is low for small values of $d$, but for $d\gtrsim 8$ the accuracy is high and becomes insensitive to the specific value of $d$. The inset in Figure 2A shows the output of UMAP algorithm [68] in two dimensions when applied to the vectors $\bm{\lambda}\in\mathbb{R}^{32}$ used in the distance $d_{\rm u}$. Each symbol corresponds to one temporal graph, the color and marker shape indicate its generative model and the marker size is proportional to the number of graph nodes. We observe that graph size does not appear to affect distances systematically and that the $\bm{\lambda}$ vectors used by the distance $d_{\rm u}$ allow us to discriminate all four generative models. Indeed, we have verified that the distances $d_{\rm u}$ between the graphs and the distances between their respective UMAP embedded vectors are strongly correlated, with a highly significant Spearman correlation coefficient of $\sim 0.92$.
We now carry out a test using the SocioPatterns empirical data sets describing time-resolved, close-range proximity interactions in schools over multiple days. These data sets are Primary school, High school 1, High school 2, High school 3. In these graphs, the nodes can be grouped according to school classes. In the primary school data set, there are ten classes, subdivided into two sections and five grades. In the high school data sets, the classes are defined by the main topics studied by the students: these topics can be maths and physics, physics and chemistry, or biology. The $2011$ data set is composed of three classes on physics and chemistry; the $2012$ data set has two additional classes on physics and maths; the $2013$ data set contains another physics and maths class and three biology classes. The metadata grouping nodes into classes are provided alongside the temporal networks data sets. We then assign one label to each class: the label is Primary for all primary school classes, while for the high school data sets, we only make the distinction between biology classes, labeled as Bio, and all others, labeled as Other.
Given a pair of classes $c_{1}$, $c_{2}$ from the same data set, we construct triplets $(c_{1},c_{2},day)$ inducing sub-temporal graphs formed by all the interactions on that day between the students of the two classes. For each sub-temporal graph, we compute the vector $\bm{\lambda}$ of Equation (4) and further decrease its dimensionality to $d=2$ using UMAP. We find that the Spearman correlation between the distances $d_{\rm u}$ and the distances between the UMAP embedding vectors is very high, namely equals $0.88$. Finally, we plot the resulting embedding in Fig. 2B, in which each point corresponds to a triplet. The color and marker coding denote the types of the two interacting classes. The plot shows that the distance can disentangle the networks formed from primary school data from those containing high school data. Moreover, a less neat but still visible distinction exists in the embeddings of (Bio-Other) with respect to the other triplets, likely because biology students spent more time in the labs and had different interaction patterns. This result confirms the ability of our distance to capture information on the temporal graph at multiple levels.
@@
One required property for matched distance is to distinguish between graphs that differ only upon a node re-labeling. We work with the synthetic temporal graphs defined in section Discriminating between classes of temporal graphs to evaluate our distance in accomplishing this task. For each random graph model, we i) generate an instance of the random graph with $n=1000$ nodes; ii) randomly select a fraction $\alpha$ of nodes and shuffle their labels; iii) compute the distance between the original graph and the partially re-labeled one. This is repeated for $25$ different realizations of the partial re-labeling. Figure 2C shows the distance $d_{\rm m}$ (rescaled by the graph size) as a function of $\alpha$. The plots confirm for all the considered generative models that the distance depends on partial node re-labeling, with a positive distance as soon as $\alpha>0$, and increases with the fraction $\alpha$ of relabeled nodes. Notably, the plot suggests a high discriminative power even for small $\alpha$ values, as the distance increases rapidly with $\alpha$.

### v3

@@ Introduction
Graphs are ubiquitous mathematical entities formed by a set of nodes and one of edges connecting pairs of nodes [1, 2, 3]. They can model a wide range of interacting systems, such as social [4, 5], technological [6], spatial [7] and biological networks [8], and owe their popularity to their ability to encode the complex relational structure of these systems. Real-world systems, moreover, often have temporal properties that cannot be encoded in static graphs and call for modeling based on time-resolved network representations, known as temporal graphs [9]. Examples of these include transportation [10] and ecological networks [11], human close-range interactions [12, 13, 14, 15, 16], collaboration networks [17, 18, 19], etc.
@@
In the case of temporal graphs, the temporal dimension adds to the complexity of graph comparison. Many of the distance measures mentioned above for static graphs were used to study temporal graphs [35], typically to compare the state of the graph at different points in time and to identify anomalies and change points in the temporal evolution of the graph [36, 37, 38, 39, 40]. Here instead, we study the problem of comparing temporal graphs, each taken as a whole. Unlike the snapshot-based methods mentioned above [36, 37, 38, 39, 40], we rely on a temporal embedding with a dimensionality that is independent of the temporal span of the graphs, so that we can compare temporal graphs with arbitrary, and different, temporal spans. Few notions of distance were proposed to compare temporal graphs with arbitrary time spans. In particular, the method of Ref. [41] is based on a set of features that can be tailored to a specific research question (similarly to [29] for static graphs). Ref. [42] on the other hand addresses the same problem of temporal graph comparison we tackle here, but it assumes the graphs to be matched and it is computationally inefficient (the graph dissimilarity is obtained solving an NP-hard optimization problem and the authors propose an approximation running in $\mathcal{O}(n^{3})$ operations, where $n$ is the number of graph nodes). Finally, Ref. [43] only considers the unmatched case and focuses on a specific property of temporal graphs, namely the comparison of temporal paths’ statistics, to provide a definition of dissimilarity.
Here, we propose a computationally efficient method to compute an actual distance metric between pairs of temporal graphs, considering both the matched and unmatched cases. We build our comparison on top of graph embeddings, leveraging ideas from the local and global approaches mentioned above and encoding differences at the topological and temporal levels. Commented Python code that implements the approach described here is publicly available at the link provided in Ref. [44].
Results
@@
A temporal graph $\mathcal{G}(\mathcal{V},\mathcal{E},\mathcal{W})$ is a tuple $(\mathcal{V},\mathcal{E},\mathcal{W})$, where $\mathcal{V}$ is a set of $n$ nodes, $\mathcal{E}$ a set of temporal edges between pairs of nodes, and $\mathcal{W}$ a set of edge weights. For a discrete variable $t$ denoting time, all temporal edges $e\in\mathcal{E}$ have the form $e=(i,j,t)$, representing an interaction between nodes $i$ and $j$ at time $t$ and $w_{e}$ indicates the positive weight of edge $e$. We say an edge $(ij)$ is active at time $t$ if $(i,j,t)\in\mathcal{E}$. A node $i$ is active at time $t$ if it belongs to an active edge at $t$.
@@
Figure 1: Distance computation between two temporal graphs. The inputs to the distance function are two temporal graphs $\mathcal{G}_{1}$ (top, orange nodes) and $\mathcal{G}_{2}$ (bottom, green nodes), generally with a different number of snapshots ($T_{1}=5$, $T_{2}=4$, in the example) and a different number of nodes $n$ (here $n_{1}=6$, $n_{2}=5$). Each graph is represented as a matrix $P\in\mathbb{R}^{n\times n}$, according to Eq. (1), with entry $(ij)$ encoding the probability that a random walker goes from $i$ to $j$ following a time-respecting path (the walker’s position at each time point is indicated in blue). The matrices $P_{1},P_{2}$ are then embedded using the EDRep algorithm, mapping them to $X_{1}\in\mathbb{R}^{n_{1}\times d}$ and $X_{2}=\mathbb{R}^{n_{2}\times d}$. Finally, the matched – Eq. (5) – and unmatched – Eq. (6) – distances are calculated. Notice that a necessary condition to compute $d_{\rm m}$ is that $n_{1}=n_{2}$.
In the following, we define two types of distance functions, $d_{\rm m}$, $d_{\rm u}$ between matched and unmatched graphs, respectively. The inputs to these functions are two temporal graphs, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, with potentially different numbers of time points (snapshots) $T_{1}$ and $T_{2}$. In both cases, the distances should fulfill the following standard requirements [46]: i) non-negativity; ii) separation axiom; iii) symmetry; iv) triangle inequality. In addition, both distances should be able by design to detect differences induced by the permutation of time indices. The matched distance should also differentiate between temporal networks differing only by a permutation of node indices, while the unmatched one should be invariant under such transformation.
Our proposed distances are computed in two steps: i) we generate a temporal graph embedding, given by a matrix of size $(n\times d)$ (with $d$ independent of $T$), that encodes relevant topological and temporal properties by leveraging time-respecting random walks on the temporal network; ii) we define the distance in the embedded space, treating separately the matched and the unmatched cases. The distance computation pipeline is illustrated in Fig. 1. The steps outlined above are described in detail in the following two Sections titled Temporal graph embedding and Defining an embedding-based distance.
@@
Graphs are rich data representations and it is challenging to define simple mathematical operations to manipulate them. Graph embeddings provide an interesting avenue for doing so, as they yield a representation in vector space that preserves some of the graphs’ relevant properties [47, 48, 49, 50]. In the case of temporal graphs, the additional challenge of dealing with their temporal dimension has been tackled in many different ways: Ref. [51] represents time as an additional dimension of an adjacency tensor; in evolutionary spectral clustering [52, 53, 54, 55, 56], a different embedding is obtained for each snapshot, and smoothness conditions are implemented on their temporal evolution; more recent work [57, 58, 59] introduced approaches based on time-respecting random walks, which encode temporal properties relevant for dynamical processes occurring on temporal graphs; Refs. [39, 40] also build one embedding per snapshot for studying the graph evolution and comparing different time points; the authors argue that a graph comparison metric should depend on the whole graph structure and thus define graph instead of node embeddings.
Following an approach similar to Ref. [59], we start from time-respecting random walks to generate temporal graph embeddings, which we use to define a distance based on global temporal graph properties. The embeddings are obtained as the solution to an optimization problem. We define a loss function whose argument is a transition matrix $P$ that is the limiting distribution of time-respecting random walks over the temporal graph $\mathcal{G}$ to be embedded. $T$ being the number of snapshots of $\mathcal{G}$, we consider an ensemble of random walks of length $\ell$ drawn uniformly at random between $1$ and $T$, and starting from a randomly chosen node at time $t=T-\ell+1$; at each time step, the walker, situated, e.g., on node $i$, can either move to a random neighbor of $i$ in the snapshot at time $t$ or stay in place on $i$; if the walker is situated on a node that is not active at time $t$, i.e., which has no neighbors in this snapshot, the walker stays in place with probability one. The graph snapshots on the left of Fig. 1 show examples of such time-respecting random walks on the input graphs.
Specifically, given a temporal graph $\mathcal{G}$, with $W_{t}$ the snapshot adjacency matrices at times $t=1,\cdots,T$, we define for each snapshot $\hat{L}_{t}=(D_{t}+I_{n})^{-1}(W_{t}+I_{n})$, where $D_{t}$ is the degree matrix $D_{t}={\rm diag}(W_{t}\mathbf{1}_{n})$ of the snapshot, $I_{n}$ is the identity matrix of size $n\times n$, and $\mathbf{1}_{n}$ is the $n$-sized unity vector. We then define the matrix $P$ as
@@
The matrices $\hat{L}_{t}$ and $P$ are sometimes referred to as temporal transition matrix and global transition matrix, respectively. Time-respecting random walks depend on several properties of the temporal graph, that may include, for instance, the presence of a non-Markovian behavior, and of a broad distribution of the edge activity and inter-event durations [60]. The limiting probability matrix $P$ of time respecting random walks encodes the aforementioned properties and it is sensitive to the permutation of time indices. We remark that the random walks we use can be of arbitrary length. Hence, the matrix $P$ is, in principle, sensitive to all time scales of the temporal graph at hand.
Given the above probability matrix $P$ and a chosen embedding dimensionality $d$, we now generate, for each node $i$ of $\mathcal{G}$, a unitary node embedding vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$. To this end, we define $X\in\mathbb{R}^{n\times d}$ as a matrix where each row represents an embedding vector. We then express the loss function as:
$\mathcal{L}_{P}(X)$ $=-\sum_{i,j\in\mathcal{V}}\left[P_{ij}~{}{\rm log}~{}Q_{ij}(X)-\frac{\mathbf{x}_{i}^{T}\mathbf{x}_{j}}{n}\right]\,\,,$ (2)
$Q_{ij}(X)$ $=\frac{e^{\mathbf{x}_{i}^{T}\mathbf{x}_{j}}}{\sum_{k\in\mathcal{V}}e^{\mathbf{x}_{i}^{T}\mathbf{x}_{k}}}:=\frac{e^{\mathbf{x}_{i}^{T}\mathbf{x}_{j}}}{Z_{i}}\,\,.$ (3)
@@
We want to define a distance between temporal graph embeddings satisfying the requirements stated above. Notice that the embedding vectors $\mathbf{x}$ are defined up to an orthogonal transformation applied to the rows of $X$. In fact, suppose $R\in\mathbb{R}^{d\times d}$ is an orthogonal matrix, and $\tilde{\mathbf{x}}_{i}=R\mathbf{x}_{i}$, then $\mathbf{x}_{i}^{T}\mathbf{x}_{j}=\mathbf{\tilde{x}}_{i}^{T}\mathbf{\tilde{x}}_{j}$ and the loss function of Eq. (2) is $\mathcal{L}_{P}(\tilde{X})=\mathcal{L}_{P}(X)$. Hence, our distance must be invariant under orthogonal transformations. Moreover, as discussed above, the distance for unmatched graphs must also be invariant with respect to node permutations.
Given these requirements, we now discuss separately our choices for the the matched and unmatched cases. In the Evaluation with empirical data section, we show the ability of our distances to discriminate between temporal graph classes that differ for topological or temporal properties.
Distance definition for matched graphs
@@
We can easily verify that $d_{\rm m}$ is a distance, as it satisfies all the requirements mentioned above and is generally not invariant under the permutation of node indices. Notice that, even if $d_{\rm m}$ can be computed for any pair of graphs with the same size, this distance is only meaningful for matched graphs (i.e., with known matching). In Eq. (5) indeed, the term $X_{1}^{T}X_{2}$ implies that $X_{1}$ and $X_{2}$ have the same size and that their rows are ordered according to the same node indices.
Distance definition for unmatched graphs
@@
In the Methods section we show that the covariance matrices $X_{1}^{T}X_{1}$ and $X_{2}^{T}X_{2}$ (of size $d\times d$) are indeed an appropriate argument of the distance function $d_{\rm u}$, since they are independent of the ordering of the nodes and any other orthogonal transformation applied to the columns of $X_{1}$ and $X_{2}$. However, a direct comparison of the entries of $X_{1}^{T}X_{1}$ and $X_{2}^{T}X_{2}$ would require a one-to-one correspondence between the embedding dimensions, i.e., the columns of $X_{1}$ and $X_{2}$. This mapping is unavailable because the EDRep algorithm, like most vector-space embedding techniques, makes no guarantees on the correspondence between the dimensions of the embedding vectors related to different graphs. To tackle this issue, we thus compare the spectra, which are invariant with respect to orthogonal transformations of the rows of the embedding matrices. This choice takes inspiration from spectral distance definitions for static graphs, such as Ref. [33], in which graphs are compared by computing the distance between the ordered sets of the eigenvalues of their matrix representations (e.g., adjacency, Laplacian, non-backtracking, etc.)
@@
As discussed in the Methods, the computational complexity of the embedding step is $\mathcal{O}\left(nd^{2}+d\cdot{\rm min}(n^{2},E)\right)$, where $E=\sum_{t}|\mathcal{E}_{t}|$ is the total number of temporal edges, $n$ is the number of graph nodes and $d$ is the embedding dimension. On top of this, the distance $d_{\rm m}$ is then obtained with $\mathcal{O}(nd^{2}+d^{2})$ additional operations, while $d_{\rm u}$ requires $\mathcal{O}(nd^{2}+d^{3})$ additional operations.
Evaluation with empirical data
@@
Figure 2: Validation of the distances on graphs of varying size. Panel A: Accuracy of the distance-based clustering against the ground truth classes in terms of normalized mutual information (NMI), as a function of the embedding dimension $d$ used in the embedding step shown in Figure 1. The clustering task consists in recognizing four classes of synthetic temporal graphs with an unsupervised algorithm based on the distance $d_{\rm u}$. The classes are obtained by generating a graph from either of four models (stochastic block model (SBM), configuration model (CM), Erdős-Renyi (ER), and geometric model (GM)) with constant degree equal to $4.8$, and the temporal component is obtained by sampling the edge activity of an empirical graph, as detailed in the main text. Inset of panel A: Scatter plot of UMAP dimensionality reduction in two dimensions of the vector $\bm{\lambda}$ appearing in the definition of $d_{\rm u}$ given in Eq. (6), with $d=32$. Each point refers to a temporal graph; the color and marker style refer to the generative model of its static component, while the marker size is proportional to the number of nodes. Panel B: $2$-dimensional UMAP embedding of $\bm{\lambda}$ for the temporal graph obtained selecting two classes and a day of interaction for all possible $(c_{1},c_{2},\mbox{Day})$ triplets in the SocioPatterns datasets describing temporal graphs of human proximity in schools. Each point refers to a triplet and the color and marker style are assigned according to the class labels: the primary school classes form a group on their own and the other three groups (Other-Other, Bio-Other, Bio-Bio) belong to the high school datasets, where Bio are the biology classes, and Other the remaining ones. In all cases, the temporal networks are aggregated over a scale of $t_{\rm res}=10$ minutes.
@@
Figure 2A shows the normalized mutual information (NMI) between the known temporal graph classes (associated with the four generative models) and the inferred cluster labels, as a function of the embedding dimension $d$. The discrimination performance is low for small values of $d$, but for $d\gtrsim 8$ the accuracy is high and becomes insensitive to the specific value of $d$. The inset in Figure 2A shows the UMAP [69] $2$-dimensional embedding of the vectors $\bm{\lambda}\in\mathbb{R}^{32}$ used in the distance $d_{\rm u}$. Each symbol corresponds to one temporal graph, the color and marker shape indicate its generative model and the marker size is proportional to the number of graph nodes. We observe that the graph size does not appear to affect distances systematically and that the $\bm{\lambda}$ vectors used by the distance $d_{\rm u}$ allow us to discriminate between all four generative models. Indeed, we verified that the distances $d_{\rm u}$ between graph pairs and the Euclidean distances between their respective UMAP embeddings are strongly correlated, with a highly significant Spearman correlation coefficient of $\sim 0.92$.
We now carry out a test using the distance $d_{\rm u}$ on empirical temporal graphs. We look at the challenging setting in which the graph pairs considered have comparable sizes and time spans and are collected in similar contexts. To do so, we use the SocioPatterns datasets collected in schools Primary school, High school 1, High school 2, High school 3, described in Table 1, representing time-resolved, close-range proximity interactions of individuals grouped in multiple school classes, over several days. We extract sub-temporal graphs from these datasets and compare them with one another using the distance $d_{\rm u}$. We associate the school classes with a label: Primary for all primary school classes; Bio for the high school classes on biology; Other for the remaining classes.
We consider a pair of classes $(c_{1},c_{2})$ from the same dataset, and we build a sub-graph restricted to all the interactions on a given day “Day” between the students of the classes $c_{1}$ and $c_{2}$. We finally compute the $2$-dimensional UMAP embedding from the vector $\bm{\lambda}$ appearing in Eq. (6). Figure 2B shows the result of this procedure for all possible triplets $(c_{1},c_{2},\mbox{Day})$. Class labels are indicated by the markers’ color and style. We observe a very high correlation between the distances $d_{\rm u}$ of graph pairs and the Euclidean distances between their UMAP embedding vectors, with a Spearman correlation of 0.880.88~{}0.880.88. That is, our distance definition can tell apart the networks extracted from the primary school dataset from those coming from the high school dataset. Moreover, a weaker but still visible separation exists in the embeddings of (Bio-Other) with respect to the other triplets. This is likely because biology classes were often held in school laboratories rather than in classrooms, leading to different interaction patterns. Our distance can discriminate between temporal networks that are, in principle, similar (school classes over one day) with no major structural or temporal differences.
Figure 3: Detection of partial node relabeling. Normalized matched distance $d_{\rm m}/n$ between a temporal graph and itself, upon partial node re-labeling, as a function of the fraction $\alpha$ of re-labeled nodes. The plots refer to the four generative models used in Figure 2A: stochastic block model (SBM), Erdős-Renyi (ER), configuration model (CM), and geometric model (GM). with average degree equal to $4.8$. For each graph, the temporal component is obtained by sampling the edge activity of an empirical graph, as detailed in the main text. The barely visible shadow line is the standard deviation of the distance across $25$ randomizations of the partial re-labeling.
@@
One required property for matched distance is to distinguish between graphs that differ only upon a node re-labeling. To evaluate the ability of our distance to accomplish this task, we work with the synthetic temporal graphs defined in section Discriminating between classes of temporal graphs. For each random graph model, we i) generate an instance of the random graph with $n=1000$ nodes; ii) randomly select a fraction $\alpha$ of nodes and shuffle their labels; iii) compute the distance between the original graph and the partially re-labeled one. This is repeated for $25$ different realizations of the partial re-labeling. Figure 3 shows the distance $d_{\rm m}$ (rescaled by the graph size) as a function of $\alpha$. The plots confirm that the distance depends on partial node re-labeling for all the considered generative models, with a positive distance as soon as $\alpha>0$, and increases with the fraction $\alpha$ of relabeled nodes. Notably, the plot suggests a high discriminative power even for small $\alpha$ values, as the distance increases rapidly with $\alpha$.